A mercury barometer reads 744 mm on the roof of a building and 761 mm on the ground. Assuming a constant value of 1.29 kg/m3 for the density of air, determine the height of the building.
Answer #1

(rho)g(h - h') = rho'*g*H
13534*(761-744)*10^-3 = 1.29*H
H = 178.355 m